Tickets to a community theater cost $25 for adults and $5 for kids. At a recent performance, a total of 305 tickets were sold for $6025. How many tickets of each kind were sold
Answer by jorel1380(3719)

l+r=305
25l+5r=6025
5l+5r=1525
20l=4500
l=225
r=80
225 adult tickets were sold, along with 80 kids tickets..